If Al can paint a room in 20 hours, and Bill can paint a room in 30 hours, how long will it take to paint a room if they work together?
Answer by richwmiller(17219)

1/20+1/30=1/x
x=12 hours for them together
check
12/20+12/30=1
6/10+4/10=1